# Advancing Surgical VQA with Scene Graph Knowledge

### v1

@@ 3.3.2 Results on Cholec80-VQA
Table 5: Results on the Cholec80-VQA. SSG-QA-Net achieves better results than the state-of-the-art models, even w.r.t SurgicalGPT, which contains a heavy sequence decoding module of GPT-2
@@
VisualBert  (Li et al., 2019) 209.8M 89.7 62.9 63.3
VisualBert Resmlp  (Seenivasan et al., 2022) 184.7M 89.8 62.7 63.4
SurgicalGPT  (Seenivasan et al., 2023) 234.5M 87.5 57.5 57.9
SSG-QA-Net 145.3M 90.6 64.4 63.7
We also conduct the experiments on the other publicly available surgical VQA dataset Cholec80-VQA. As illustrated in Table 5, SSG-QA-Net significantly outperforms the SurgicalGPT  (Seenivasan et al., 2023), which requires heavy sequence decoding using GPT-2 architecture. This highlights that the bottleneck of the current surgical VQA problem lies in the visual scene understanding instead of text generation. Also, even using YOLOv7 for object detection, our model achieves higher performance with fewer parameters than prior works, verifying its efficiency while maintaining higher performance metrics.
@@ 4 Conclusion
In this paper, we tackle the problem of visual question answering (VQA) in the context of fine-grained surgical scene understanding. First, we introduce a new dataset called SSG-QA, which uses a surgical scene graph as an underlying representation and a question-answer generation engine to generate diverse, geometrically grounded, and surgical-action-oriented question-answer pairs. The question-answer pairs are also sampled to mitigate the question-condition bias that exists in the current surgical VQA datasets. We also propose a novel model called SSG-QA-Net to explicitly incorporate scene knowledge and object-wise local features in the VQA model design to improve the reasoning ability on complex questions. The results show that SSG-QA-Net outperforms existing baseline models by a large margin.
@@ 5 Acknowledgments
This work was supported by French state funds managed by the ANR within the National AI Chair program under Grant ANR-20-CHIA-0029-01 (ChairAI4ORSafety) and within the Investments for the future program under Grant ANR-10-IAHU-02 (IHU Strasbourg). This work was granted access to the HPC resources of IDRIS under the allocations AD011013704R1, AD011011631R2, and AD011011631R3 made by GENCI.
@@ 5.1 SSG Dataset
In this section, we first illustrate the statistics of the SSG-QA dataset and show that it is more challenging and diverse than the prior works. Then, we show the process of assigning location attributes to the surgical objects. Finally, we visualize some samples of the scene graphs that we use to generate question-answer pairs.
Table 7: Dataset statistics comparison. SSG-QA is more challenging and balanced as it includes more attributes and complexities in the questions.
Dataset EndoVis-18-VQA Cholec80-VQA SSG-QA
@@ 5.1.1 Dataset Statistics
As shown in Tab. 7, SSG-QA dataset contains the more challenging and diverse question-answer pairs. Specifically, the average length of question is $38.9$ words compared to $5$ words and $6.5$ words from EndoVis-18-VQA and Cholec80-VQA. Also, by using the question engine to automatically generate the question-answer pairs, SSG-QA contains a more diverse and different pattern of question-answer pairs for each surgical scene.
@@ 5.1.2 Location Attribute and Triplet Annotation
Fig. 5: SSG-QA Dataset - location attribute: (a) we use triplet annotations for action-related question-answer generation; (b) we include a spatial attribute for each object, e.g., top-right, bottom-left; (c) we use tool and anatomy bounding boxes to generate question-answer pairs.
@@ 5.1.3 Scene Graphs
Fig. 6: SSG-QA Dataset - scene graphs: The questions are generated by exploiting the surgical scene graph. The surgical scene graph consists of spatial attributes and action relations.
Fig. 7: SSG-QA Dataset - different complexities: We generate different type of question-answer pairs for each surgical scene image. The text in red indicates the reasoning step in the image, e.g., the fifth question requires two steps of reasoning to solve, therefore categorized into single-and type.
Fig. 8: Qualitative Results: Comparison of SSG-QA-Net with baseline surgical VQA models on the question-answer pairs based on spatial, color, counting, anatomy, and tool presence attributes.
@@ 5.1.4 Complexities of Question-answer Pairs
As the SSG-QA dataset is generated with a question engine and question templates automatically, we control the complexity of question-answer pairs by altering the question templates. The question-answer pairs with different complexities also offer the diagnostic ability to pinpoint the weakness of the VQA model. As shown in Fig. 7, we have $5$ kinds of question-answer pairs with $3$ kinds of complexities. Here, the complexity is measured based on how many reasoning steps are required to answer the question. For example, the fifth question-answer in Fig. 7 requires two steps of reasoning to solve, highlighted in red. As a result, this question is more complex than the others.
Fig. 9: Qualitative Results: Comparison of SSG-QA-Net with baseline surgical VQA models on the question-answer pairs based on surgical action triplets.
Fig. 10: Qualitative Results: It shows that the questions from our SSG-QA dataset require visual reasoning to solve. Also, our SSG-QA-Net with scene knowledge outperforms the prior baselines in terms of complex questions.
@@ 5.2 Qualitative Results
In this section, we compare our SSG-QA-Net to the state-of-the-art surgical VQA models in terms of the question that require the understanding of attributes (Sec. 5.2.1), action triplet occurrence (Sec. 5.2.2) and reasoning steps (Sec. 5.2.3) to solve.
@@ 5.2.1 Attributes
As shown in Fig. 8, the SSG-QA-Net model outperforms the prior works when the questions contain attributes, such as location and color. For example, the second subfigure in Fig. 8 shows that the SSG-QA-Net model understands the location attributes such as “top-mid” and “bottom-mid” to resolve the question. Compared to the other methods, it contains more detailed knowledge about the surgical scene.
@@ 5.2.2 Action Triplets
We also compare the SSG-QA-Net model to the prior works on the questions that require an understanding of surgical action triplet. Fig. 9 shows that the SSG-QA-Net model can answer the questions that are generated based on the triplet annotations.
@@ 5.2.3 Reasoning Steps
Fig. 10 shows that SSG-QA-Net can perform multiple steps of reasoning to generate the correct answer to complex questions. Compared to the prior works, it achieves better results when the questions are more complex.

### v3

@@ 3.3.2 Results on Cholec80-VQA
Table 5: Results on the Cholec80-VQA. SSG-VQA-Net achieves better results than the state-of-the-art models, even w.r.t SurgicalGPT, which contains a heavy sequence decoding module of GPT-2
@@
VisualBert  [15] 209.8M 89.7 62.9 63.3
VisualBert Resmlp  [20] 184.7M 89.8 62.7 63.4
SurgicalGPT  [19] 234.5M 87.5 57.5 57.9
SSG-VQA-Net 145.3M 90.6 64.4 63.7
We also conduct the experiments on the other publicly available surgical VQA dataset Cholec80-VQA. As illustrated in Table 5, SSG-VQA-Net significantly outperforms the SurgicalGPT  [19], which requires heavy sequence decoding using GPT-2 architecture. This highlights that the bottleneck of the current surgical VQA problem lies in the visual scene understanding instead of text generation. Also, even using YOLOv7 for object detection, our model achieves higher performance with fewer parameters than prior works, verifying its efficiency while maintaining higher performance metrics.
@@ 4 Conclusion
In this paper, we tackle the problem of visual question answering (VQA) in the context of fine-grained surgical scene understanding. First, we introduce a new dataset called SSG-VQA, which uses a surgical scene graph as an underlying representation and a question-answer generation engine to generate diverse, geometrically grounded, and surgical-action-oriented question-answer pairs. The question-answer pairs are also sampled to mitigate the question-condition bias that exists in the current surgical VQA datasets. We also propose a novel model called SSG-VQA-Net to explicitly incorporate scene knowledge and object-wise local features in the VQA model design to improve the reasoning ability on complex questions. The results show that SSG-VQA-Net outperforms existing baseline models by a large margin.
@@ 5 Acknowledgments
This work has received funding from the European Union (ERC, CompSURG, 101088553). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was also partially supported by French state funds managed by the ANR under Grants ANR-20-CHIA-0029-01 and ANR-10-IAHU-02. This work was granted access to the HPC resources of IDRIS under the allocation AD011013704R1 and AD011011631R2 made by GENCI. This work was granted access to the HPC resources of IDRIS under the allocations AD011013704R1, AD011011631R2, and AD011011631R3 made by GENCI. The authors would like to acknowledge the High Performance Computing Center of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. Part of the computing resources were funded by the Equipex Equip@Meso project (Programme Investissements d’Avenir) and the CPER Alsacalcul/Big Data.
@@ 5.1 SSG Dataset
In this section, we first illustrate the statistics of the SSG-VQA dataset and show that it is more challenging and diverse than the prior works. Then, we show the process of assigning location attributes to the surgical objects. Finally, we visualize some samples of the scene graphs that we use to generate question-answer pairs.
Table 7: Dataset statistics comparison. SSG-VQA is more challenging and balanced as it includes more attributes and complexities in the questions.
Dataset EndoVis-18-VQA Cholec80-VQA SSG-VQA
@@ 5.1.1 Dataset Statistics
As shown in Tab. 7, SSG-VQA dataset contains the more challenging and diverse question-answer pairs. Specifically, the average length of question is $38.9$ words compared to $5$ words and $6.5$ words from EndoVis-18-VQA and Cholec80-VQA. Also, by using the question engine to automatically generate the question-answer pairs, SSG-VQA contains a more diverse and different pattern of question-answer pairs for each surgical scene.
@@ 5.1.2 Location Attribute and Triplet Annotation
Fig. 5: SSG-VQA Dataset - location attribute: (a) we use triplet annotations for action-related question-answer generation; (b) we include a spatial attribute for each object, e.g., top-right, bottom-left; (c) we use tool and anatomy bounding boxes to generate question-answer pairs.
@@ 5.1.3 Scene Graphs
Fig. 6: SSG-VQA Dataset - scene graphs: The questions are generated by exploiting the surgical scene graph. The surgical scene graph consists of spatial attributes and action relations.
Fig. 7: SSG-VQA Dataset - different complexities: We generate different type of question-answer pairs for each surgical scene image. The text in red indicates the reasoning step in the image, e.g., the fifth question requires two steps of reasoning to solve, therefore categorized into single-and type.
Fig. 8: Qualitative Results: Comparison of SSG-VQA-Net with baseline surgical VQA models on the question-answer pairs based on spatial, color, counting, anatomy, and tool presence attributes.
@@ 5.1.4 Complexities of Question-answer Pairs
As the SSG-VQA dataset is generated with a question engine and question templates automatically, we control the complexity of question-answer pairs by altering the question templates. The question-answer pairs with different complexities also offer the diagnostic ability to pinpoint the weakness of the VQA model. As shown in Fig. 7, we have $5$ kinds of question-answer pairs with $3$ kinds of complexities. Here, the complexity is measured based on how many reasoning steps are required to answer the question. For example, the fifth question-answer in Fig. 7 requires two steps of reasoning to solve, highlighted in red. As a result, this question is more complex than the others.
Fig. 9: Qualitative Results: Comparison of SSG-VQA-Net with baseline surgical VQA models on the question-answer pairs based on surgical action triplets.
Fig. 10: Qualitative Results: It shows that the questions from our SSG-VQA dataset require visual reasoning to solve. Also, our SSG-VQA-Net with scene knowledge outperforms the prior baselines in terms of complex questions.
@@ 5.2 Qualitative Results
In this section, we compare our SSG-VQA-Net to the state-of-the-art surgical VQA models in terms of the question that require the understanding of attributes (Sec. 5.2.1), action triplet occurrence (Sec. 5.2.2) and reasoning steps (Sec. 5.2.3) to solve.
@@ 5.2.1 Attributes
As shown in Fig. 8, the SSG-VQA-Net model outperforms the prior works when the questions contain attributes, such as location and color. For example, the second subfigure in Fig. 8 shows that the SSG-VQA-Net model understands the location attributes such as “top-mid” and “bottom-mid” to resolve the question. Compared to the other methods, it contains more detailed knowledge about the surgical scene.
@@ 5.2.2 Action Triplets
We also compare the SSG-VQA-Net model to the prior works on the questions that require an understanding of surgical action triplet. Fig. 9 shows that the SSG-VQA-Net model can answer the questions that are generated based on the triplet annotations.
@@ 5.2.3 Reasoning Steps
Fig. 10 shows that SSG-VQA-Net can perform multiple steps of reasoning to generate the correct answer to complex questions. Compared to the prior works, it achieves better results when the questions are more complex.